OpenStudy (anonymous):

Factor the polynomial x^8 - 98x^4 + 1 into 2 non-constant polynomials with integral coefficients please explain what does non-constant polynomials with integral coefficients mean and explain each step you did

hero (hero):

Are you supposed to just factor it?

OpenStudy (anonymous):

not sure i want to understand what this question is exactly asking =/

OpenStudy (anonymous):

From Mathematica:\[\left(x^4-10 x^2+1\right) \left(x^4+10 x^2+1\right) \]

OpenStudy (anonymous):

thats for factoring but do you know what does non-constant polynomials with integral coefficients mean?

OpenStudy (anonymous):

Meaning the coeffcients much be an interger which they are.

OpenStudy (anonymous):

Meaning the coeffcients much be an interger which they are.

OpenStudy (anonymous):

Meaning the coeffcients much be an interger which they are.

OpenStudy (anonymous):

Meaning the coeffcients much be an interger which they are.

OpenStudy (anonymous):

Meaning the coeffcients much be an interger which they are.

OpenStudy (anonymous):

Meaning the coeffcients much be an interger which they are.

OpenStudy (anonymous):

Meaning the coeffcients much be an interger which they are.

OpenStudy (unklerhaukus):

\[x^8 - 98x^4 + 1\]let \(x^4=u\) \[u^2 - 98u + 1\] complete the square !